OpenStudy (anonymous):

Simplify the product using the rules for exponents: 19^3 × 19^4

OpenStudy (anonymous):

a^m * a^n = a^(m+n)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

here a =19 m =3 n = 4

OpenStudy (anonymous):

just subs in the equation

OpenStudy (anonymous):

Select one of the options below as your answer: A. 19 × 7 B. 1912 C. 19−1 D. 197

OpenStudy (anonymous):

it is 19^7

OpenStudy (anonymous):

Ignore Shane_B in the first post. Yahoo! has this one covered.

OpenStudy (anonymous):

haha ok

OpenStudy (shane_b):

I typoed...my bad.

OpenStudy (anonymous):

No both are wrong the correct answer is 19(3*4)

OpenStudy (shane_b):

No :)

OpenStudy (shane_b):

It should be 19^7...you only multiply the exponents if you're raising an exponent to another exponent.

OpenStudy (shane_b):

@blm: Maybe this explanation will be better. The rule is:\[\large a^x*a^y=a^{x+y}\]Which can be easily proven with an example: \[a^2*a^3=a^5=a*a*a*a*a\]